6It takes 3.00 ns for light to travel through a pane of glass (n 1.5). How long does it take for light to travel through the same thickness of water (n 1.33)? 3.00 ns →3.0.10-13 A. 3.00 ns 3.99 ns c) 4.50 ns D. 2.66 ns E. 3.38 ns 2 108 V

Test corrections help. If you can you answer that would be great! Thanks!
Answer #1

We know

Refractive index=(speed of light in vacuum/speed of light in medium)

So n1/n2=v2/v1=t1/t2

Putting values

1.5/1.33=3/t2

t2=2.66 ns